Company ABC has EBIT = € 18038, depreciation expense = € 2200, interest expense= € 4500 and the tax rate 35%.

What is the operating cash flow?
Answer #1

Operating cash flow = EBIT * (1-Tax Rate) + Depreciation

Therefore, Operating Cash Flow = 18038 * (1-0.35) + 2200 = 13925